While the temperature remains constant at 38oC, 2.5 kg of a certain ideal gas whose Cp = 1.1724 kJ/kg-K, Cv = 0.8375 kJ/kg-K rejects 300kJ of heat. If p1 = 586kPa, find: (a) p2, (b) V1 and V2.


1
Expert's answer
2021-12-06T16:43:12-0500

"\nu=\frac{m(c_p-c_V)}R=100.8~mol,"

a)

"Q=\nu RT\ln\frac{p_1}{p_2},\implies"

"p_2=p_1e^{-\frac Q{\nu RT}}=185~kPa,"

b)

"V_1=\frac{\nu RT}{p_1}=0.44~m^3,"

"V_2=\frac{\nu RT}{p_2}=1.41~m^3."